OpenStudy (anonymous):

Write an equation of a line that is parallel and perpendicular to the above equation.

OpenStudy (anonymous):

a) \[y=\frac{ 2 }{ 3 }x + \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

do you have a point that the lines must go through?

OpenStudy (anonymous):

Yeah, given points: (-2,-3) (4,1) I honestly don't know how to do this...I have to do this for 2 more problems.

OpenStudy (anonymous):

Given the points (-2,-3) (4,1) Write the equation in: a)Point slope form - y+3=\[\frac{ 2 }{ 3 }(x+2)

OpenStudy (anonymous):

(-2,-3) for parallel and (4,1)for perpendicular?

OpenStudy (anonymous):

i'm not sure...my teacher just gave me these 2 points.

OpenStudy (anonymous):

Hmm... You usually can't do these types of problems without knowing that. do you just want to assume that(-2,-3) is for parallel and (4,1) is for perpendicular?

OpenStudy (anonymous):

is this for a geometry class?

OpenStudy (anonymous):

nope. It's algebra... D:

OpenStudy (anonymous):

Sure...i'll assume that.

OpenStudy (anonymous):

oh okay. i'm doing this in geometry XD

OpenStudy (anonymous):

,I like dk how to do this...for a. I got y=2/3x+5/3 for parallel and y=3/2x for perpendicular. But wouldn't it be that b and c is the same thing?

OpenStudy (anonymous):

... anybody know how to do it?

OpenStudy (anonymous):

so for parallel lines, the slope stays the same. so a line parallel to y=2/3x+4/3 would have a slope of 2/3. also, if it went through point (-2,-3), you would do this: y=2/3x+b -3=2/3(-2)+b -3=-4/3+b -13/3=b -4 and 1/3=b so the equation would be y=2/3x-4u1/3 it has to be in point slope form, though. WAIT a sec, is this the same? Given the points (-2,-3) (4,1) Write the equation in: a)Point slope form - y+3=\[\frac{ 2 }{ 3 }(x+2). What is y+3=\[\frac{ 2 }{ 3 }(x+2)

OpenStudy (anonymous):

can you put that in an equation format please?

OpenStudy (anonymous):

i'm sorry, im a little bit confused. you still there?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I'm trying to work it out also. x3

OpenStudy (anonymous):

okay . what is this? y+3=\[\frac{ 2 }{ 3 }(x+2)

OpenStudy (anonymous):

\[y+3=\frac{ 2 }{ 3 }(x+2)\]

OpenStudy (anonymous):

like can you make it readable?

OpenStudy (anonymous):

what'do you mean by readable? x3

OpenStudy (anonymous):

you already did it jk :)

OpenStudy (anonymous):

oh okay. :3

OpenStudy (anonymous):

so is the problem y=2/3x+4/3 like you said in the first post, or y+3=2/3(x+2)?

OpenStudy (anonymous):

y=2/3x+5/3... was the answer for parallel..the other one was the equation...

OpenStudy (anonymous):

And for perpendicular, it's y=3/2x...

OpenStudy (anonymous):

um...here a)y+3=2/3(x+2) b)y=2/3x-5/3 c)2x-3y=5

OpenStudy (anonymous):

...

OpenStudy (anonymous):

sorry working on it.

OpenStudy (anonymous):

mk... x3

OpenStudy (anonymous):

Thanks a lot btw.. :)

OpenStudy (anonymous):

this is a. i m really sorry, i have a load of homework myself. I'll keep this page up in case you need help on B and C, but ive gotta work on mine too XD and you're welcome!